### v1

@@ 1 Introduction
In recent decades, numerous exotic hadrons beyond the conventional meson and baryon schemes in the simple quark model have been observed experimentally. For some of these states, the multiquark content is manifestly revealed by their quantum numbers or masses, such as the doubly charmed tetraquark $T_{cc}(3875)^{+}\,(cc\bar{u}\bar{d})$ [1, 2], charm-strange tetraquarks $T^{*}_{\bar{c}\bar{s}0}(2870)^{0},T^{*}_{\bar{c}\bar{s}1}(2900)^{0}\,(\bar{c}\bar{s}ud)$ [3, 4], and fully charmed tetraquarks $(cc\bar{c}\bar{c})$ [5, 6, 7]. Investigating the binding mechanisms and clustering behaviors of these tetraquark states may help advance our understanding of the nonperturbative quantum chromodynamics (QCD).
Extending the quark potential model to multiquark configurations may serve as an approach to understanding the exotic hadrons in a unfied framework. The potential parameters used to calculate the tetraquark spectrum can be determined from the known meson or baryon spectra. Since no extra fits to the experimental tetraquark results are needed, the quark model can provide forward-looking predictions of various tetraquark systems. Additionally, the quark model makes no a priori assumptions about the clustering of quarks. Therefore, it can consistently describe tetraquark states with different configurations, including meson molecules and compact tetraquarks. We utilize effective few-body methods, including the Gaussian expansion method [8] and complex scaling method [9, 10, 11], to solve the Schrödinger equation and identify bound and resonant states. We further identify the spatial structures of tetraquark states by calculating their root mean square radii. We apply the framework to perform systematical studies of heavy flavor tetraquark systems, where candidates for a series of experimental states are found and predictions for future states are provided.
@@ 2.1 Hamiltonian
where $m_{i}$ and $p_{i}$ are the mass and momentum of (anti)quark $i$, $V_{ij}$ denotes the two-body interaction, and the center-of-mass kinetic energy $T_{C.M.}$ is subtracted. The interaction $V_{ij}$ takes various forms in different models. In this work, we adopt three models which contain one-gluon-exchange and quark confinement terms, including the AL1 and AP1 model proposed in Ref. [12] and the BGS model proposed in Ref. [13]. In the S-wave systems, the potential is written as
@@
Here, $\boldsymbol{\lambda}_{i}$ (replaced by $\boldsymbol{\lambda}_{i}^{*}$ for antiquark) and $\boldsymbol{S}_{i}$ are the SU(3) color Gell-Mann matrix and spin operator of (anti)quark $i$. The AL1 and BGS model contains a linear confinment term ($p=1$), while the AP1 model contains a $p=\frac{2}{3}$-power confinement term. The parameters in the models were all fitted by the meson spectra and can be found in Refs. [12, 13]. The theoretical masses of the mesons are compared with the experimental ones in Table 1. Results from all three models agree with the experimental values within tens of MeV. We expect the model uncertainties of tetraquark states to be of the same order.
Table 1: Meson mass spectra (in MeV) from the AL1, AP1 and BGS models. The experimental values taken from Ref. [14] are listed for comparison.
@@ 2.2 Complex Scaling Method
Resonant states are poles of the S matrix on the complex energy plane. The resonant wave functions are not square integrable, causing numerical difficulties for obtaining resonances. The complex scaling method [9, 10, 11] enables solving for tetraquark bound states and resonant states simultaneouly by carrying out analytical continuation of the Hamiltonian. The coordinate $\mathbf{r}$ and its conjugate momentum $\mathbf{p}$ are transformed as
@@ 2.3 Gaussian Expansion Method
The Gaussian expansion method is shown to be an effective numerical approach for solving few-body Schrödinger equation [8]. The four-body spatial wave function is expanded by the Gaussian basis functions,
@@
The incorporation of dimeson and diquark-antidiquark configurations enables descriptions of both molecular and compact states. The Gaussian basis functions can well describe both short-range and long-range correlations [8]. In addition, the matrix elements of Gaussian bases can be solved analytically [8, 15], making large-scale basis expansion calculations numerically feasible. The complete four-body wave function can be written as,
$\Psi^{J}(\theta)=\mathcal{A}\sum_{\rm{jac}}\sum_{\alpha,n_{i}}\,C^{(\rm{jac})}_{\alpha,n_{i}}(\theta)\left[\chi_{\alpha}^{S}\otimes\Phi^{L,(\rm{jac})}_{n_{1},n_{2},n_{3}}\right]^{J},$ (7)
where $\mathcal{A}$ and $C$ are the antisymmetric operator of identical particles and the expanded coefficients, respectively. $\chi_{\alpha}^{S}$ is the color-spin wave function, given by
@@
for all possible spin combinations $s_{1},s_{2},S$.
@@ 3.1 Fully heavy tetraquark
We study the S-wave fully charmed tetraquark systems using three different potential models, including the AL1, AP1 and BGS model. We find a series of resonant states, whose energies are listed in Table 2 [16]. The three models give qualitatively consistent energy spectra: Most resonances exist in all three models with similar masses and widths. Moreover, the tetraquark spectra in systems with different quantum numbers $J^{PC}$ demonstrates analogous characteristics. A lower state at $M\approx 7$ GeV with width $\Gamma\approx 75$ MeV, and a higher state at $M\approx 7.2$ GeV with width $\Gamma\approx 50$ MeV are found in the $0^{++},1^{+-},2^{++}$ systems. The lower $0^{++}$ and $2^{++}$ states are good candidates for the experimental $X(6900)$ state [5, 6, 7], while the higher $0^{++}$ and $2^{++}$ states are good candidates for the experimental $X(7200)$ state [6, 7]. Both the masses and widths agree with the experimental values within tens of MeV. The similar pattern of spectra with different $J^{PC}$ is expected, as the spin splitting term in Eq. (2) is suppressed by the heavy quark mass. However, recent analyses from the CMS collaborations suggest that the experimental states have $J^{PC}=2^{++}$ [17]. Further studies are needed to better understand and resolve the discrepancy.
Table 2: The complex energies $E=M-i\Gamma/2$ (in $\mathrm{MeV}$) of the fully charmed tetraquark resonances from three potential models. The ”?” indicates that potential signs of resonant states are observed but the pole positions cannot be obtained accurately in the present calculations. The ”-” suggests that no corresponding resonance is found.
$J^{PC}$ AL1 AP1 BGS
$0^{++}$ 698035i6980-35i 697836i6978-36i 703036i7030-36i
@@
We further investigate the S-wave fully bottomed tetraquark [16] and fully heavy tetraquark with different flavors [18] within a unified framework. A series of resonant states are predicted in $(19.7,20.0)$ GeV for the $bb\bar{b}\bar{b}$ system and $(13.2,13.5)$ GeV for the $bc\bar{b}\bar{c}$ system, which may be searched for in future experiments. From the root-mean-square (rms) radii, we find that all fully heavy tetraquark resonances exhibit a compact tetraquark configuration, with four (anti)quarks confined within the typical range of color confinement $\Lambda_{\text{Q}CD}^{-1}\sim 1$ fm. This is expected as the interactions between heavy (anti)quarks should be dominated by short-range gluon exchange. The fully strange tetraquarks, as cousins of the fully heavy tetraquarks, are also studied in Ref. [19].
@@ 3.2 Doubly heavy tetraquark
As the three potential models are found to give similar results in the fully charmed tetraquark systems, we only choose the AL1 model to investigate the S-wave doubly heavy tetraquark systems $QQ^{(\prime)}\bar{q}\bar{q}$ ($Q^{\prime}=b,c$ and q=u,dq=u,d[20]. We observe rich spatial configurations of various doubly heavy tetraquark states by calculating their rms radii. In addition to the general classification of tetraquarks into meson molecules and compact tetraquarks, we further categorize the compact tetraquark configuration into three subtypes, as illustrated in Fig. 2. In compact even tetraquark, the four (anti)quarks are evenly distributed, where their relative distances are of similar size. The two light antiquarks are shared by two heavy quarks, resembling the hydrogen atom, where two electrons are shared by two protons. In compact diquark-antidiquark tetraquark, diquark and antidiquark clusters are formed within the compact structure. In compact diquark-centered tetraquark, two heavy quarks form a compact diquark core, whereas two light antiquarks orbit around the diquark. It corresponds to the QCD helium atom, where two electrons orbit around the helium nucleus. These classifications unravel different internal structures and binding mechanisms of various tetraquark states.
@@
Table 3: The complex energies $E=M-i\Gamma/2$ (in MeV), proportions of color configurations and spatial configurations of the doubly heavy tetraquark bound and resonant states. The binding energies $\Delta E$ of the bound states are shown in the fourth column. The spatial configurations are shown in the last column, where C.E., C.DA., C.DC. and M. respectively indicate compact even tetraquark, compact diquark-antidiquark tetraquark, compact diquark-centered tetraquark and meson molecule. The constituent mesons are shown in the parentheses. The ”?” indicates that the rms radii are numerically unstable in the complex scaling transformation and the configuration of the state is undetermined.
@@ 3.3 Singly heavy tetraquark
We further study the S-wave charm-strange tetraquark ($cs\bar{q}\bar{q}$) and bottom-strange tetraquark ($bs\bar{q}\bar{q}$) using the AL1 model [21]. Table 4 lists the low-lying spectra of the two systems. We identify a series of molecular bound and resonant states, including two scalar bound states ($D\bar{K}$ and $\bar{B}\bar{K}$), two scalar resonances ($D^{*}\bar{K}^{*}$ and $\bar{B}^{*}\bar{K}^{*}$), and one vector bound state ($\bar{B}^{*}\bar{K}$). The $D^{*}\bar{K}^{*}$ molecular resonance stands as a great candidate for the experimental $T^{*}_{\bar{c}\bar{s}0}(2870)$, whose mass and width are $2872\pm 16$ MeV and $67\pm 24$ MeV [14]. The theoretical width is expected to be smaller than the experimental value due to the neglection of the width of $\bar{K}^{*}$ in the calculations.

### v2

@@ 1 Introduction
In recent decades, numerous exotic hadrons beyond the conventional meson and baryon schemes in the simple quark model have been observed experimentally. For some of these states, the multiquark content is manifestly revealed by their quantum numbers or masses, such as the doubly charmed tetraquark $T_{cc}(3875)^{+}\,(cc\bar{u}\bar{d})$ LHCb:2021vvq , LHCb:2021auc , charm-strange tetraquarks $T^{*}_{\bar{c}\bar{s}0}(2870)^{0},T^{*}_{\bar{c}\bar{s}1}(2900)^{0}\,(\bar{c}\bar{s}ud)$ LHCb:2020bls , LHCb:2020pxc , and fully charmed tetraquarks $(cc\bar{c}\bar{c})$ LHCb:2020bwg , ATLAS:2023bft , CMS:2023owd . Investigating the binding mechanisms and clustering behaviors of these tetraquark states may help advance our understanding of the nonperturbative quantum chromodynamics (QCD).
Extending the quark potential model to multiquark configurations may serve as an approach to understanding the exotic hadrons in a unfied framework. The potential parameters used to calculate the tetraquark spectrum can be determined from the known meson or baryon spectra. Since no extra fits to the experimental tetraquark results are needed, the quark model can provide forward-looking predictions of various tetraquark systems. Additionally, the quark model makes no a priori assumptions about the clustering of quarks. Therefore, it can consistently describe tetraquark states with different configurations, including meson molecules and compact tetraquarks. Previous works used the quark potential model to study doubly heavy tetraquark bound states Meng:2020knc , Deng:2021gnb . Doubly heavy tetraquark resonances were also explored using the real scaling method Meng:2021yjr , Meng:2023for . Ref. Wang:2022yes studied fully charmed tetraquark resonances using the complex scaling method. In this work, we perform systematical studies of various heavy flavor tetraquark systems, including fully heavy tetraquark $(cc\bar{c}\bar{c},bb\bar{b}\bar{b},bc\bar{b}\bar{c})$, doubly heavy tetraquark $(cc\bar{q}\bar{q},bb\bar{q}\bar{q},bc\bar{q}\bar{q})$ and singly heavy tetraquark $(cs\bar{q}\bar{q},bs\bar{q}\bar{q})$, where $q=u,d$. We utilize effective few-body methods, including the Gaussian expansion method HIYAMA2003223 and complex scaling method Aguilar1971 , Balslev1971 , Aoyama2006 , to solve the Schrödinger equation and identify bound and resonant states. We further identify the spatial structures of tetraquark states by calculating their root-mean-square radii. We find candidates for a series of experimental states and provide predictions for future states.
@@ 2.1 Hamiltonian
where $m_{i}$ and $p_{i}$ are the mass and momentum of (anti)quark $i$, $V_{ij}$ denotes the two-body interaction, and the center-of-mass kinetic energy $T_{C.M.}$ is subtracted. The interaction $V_{ij}$ takes various forms in different models. In this work, we adopt three models which contain one-gluon-exchange and quark confinement terms, including the AL1 and AP1 model proposed in Ref. SilvestreBrac1996 and the BGS model proposed in Ref. Barnes:2005pb . In the S-wave systems, the potential is written as
@@
Here, $\boldsymbol{\lambda}_{i}$ (replaced by $\boldsymbol{\lambda}_{i}^{*}$ for antiquark) and $\boldsymbol{S}_{i}$ are the SU(3) color Gell-Mann matrix and spin operator of (anti)quark $i$. The AL1 and BGS model contains a linear confinment term ($p=1$), while the AP1 model contains a $p=\frac{2}{3}$-power confinement term. The parameters in the models were all fitted by the meson spectra and can be found in Refs. SilvestreBrac1996 , Barnes:2005pb . The theoretical masses of the mesons are compared with the experimental ones in Table 1. Results from all three models agree with the experimental values within tens of MeV. We expect the model uncertainties of tetraquark states to be of the same order.
Table 1: Meson mass spectra (in MeV) from the AL1, AP1 and BGS models. The experimental values taken from Ref. ParticleDataGroup:2024cfk are listed for comparison.
@@ 2.2 Complex Scaling Method
Resonant states are poles of the S matrix on the complex energy plane. The resonant wave functions are not square integrable, causing numerical difficulties for obtaining resonances. The complex scaling method Aguilar1971 , Balslev1971 , Aoyama2006 enables solving for tetraquark bound states and resonant states simultaneouly by carrying out analytical continuation of the Hamiltonian. The coordinate $\mathbf{r}$ and its conjugate momentum $\mathbf{p}$ are transformed as
@@ 2.3 Gaussian Expansion Method
The Gaussian expansion method is shown to be an effective numerical approach for solving few-body Schrödinger equation HIYAMA2003223 . The four-body spatial wave function is expanded by the Gaussian basis functions,
@@
The incorporation of dimeson and diquark-antidiquark configurations enables descriptions of both molecular and compact states. The Gaussian basis functions can well describe both short-range and long-range correlations HIYAMA2003223 . In addition, the matrix elements of Gaussian bases can be solved analytically HIYAMA2003223 , Wu:2024ocq , making large-scale basis expansion calculations numerically feasible. The complete four-body wave function can be written as,
$\Psi^{J}(\theta)=\mathcal{A}\sum_{\rm{(jac)}}\sum_{\alpha,n_{i}}\,C^{(\rm{jac})}_{\alpha,n_{i}}(\theta)\left[\chi_{\alpha}^{S}\otimes\Phi^{L,(\rm{jac})}_{n_{1},n_{2},n_{3}}\right]^{J},$ (7)
where $J$ is the total angular momentum, $\mathcal{A}$ is the antisymmetric operator of identical particles, $C$ is the expanded coefficients. $\rm{(jac)},\alpha,$ and $n_{i}$ sum over the three types of Jacobian configurations in Fig. 1, the color-spin configurations, and the radial indices of Gaussian bases, respectively. $\chi_{\alpha}^{S}$ is the color-spin wave function, given by
@@
for all possible spin combinations $s_{1},s_{2},S$. It is worth noting that the diquark-antidiquark type color-spin basis in Eq. (8) employed in this work is equivalent to the dimeson type color-spin basis,
$\chi^{S}_{1_{c}\otimes 1_{c},s_{1},s_{2}}=\left[\left(Q_{1}\bar{Q}_{3}\right)^{s_{1}}_{1_{c}}\left(Q_{2}\bar{Q}_{4}\right)^{s_{2}}_{1_{c}}\right]^{S}_{1_{c}},$ (9)
$\chi^{S}_{8_{c}\otimes 8_{c},s_{1},s_{2}}=\left[\left(Q_{1}\bar{Q}_{3}\right)^{s_{1}}_{8_{c}}\left(Q_{2}\bar{Q}_{4}\right)^{s_{2}}_{8_{c}}\right]^{S}_{1_{c}}.$
They are both complete bases of the tetraquark color-spin degrees of freedom, and can be related to each other by unitary transformation Wu:2024euj .
2.4 Spatial structures
To investigate the clustering behaviors of tetraquark states, we use the inter-quark root-mean-square (rms) radii to distinguish between meson molecules and compact tetraquarks. In a meson molecule, the four (anti)quarks cluster into two color-singlet mesons, with their relative distances being larger than the typical color confinement scale $\Lambda_{QCD}^{-1}\sim 1\,\text{fm}$. In a compact tetraquark, the four (anti)quarks are confined together within the scale of $\Lambda_{QCD}^{-1}$. The inter-quark rms radii can be used to reveal different spatial structures. However, in the presence of identical particles, the conventional rms radii calculated using the complete wave function can be ambiguous and fail to reveal molecular configuration due to the antisymmetrization  Chen:2023syh . We propose to calculate the rms radii as
$r_{ij}\equiv\mathrm{Re}\left[\sqrt{\frac{\langle\Psi_{\mathrm{nA}}(\theta)|r_{ij}^{2}e^{2i\theta}|\Psi_{\mathrm{nA}}(\theta)\rangle}{\langle\Psi_{\mathrm{nA}}(\theta)|\Psi_{\mathrm{nA}}(\theta)\rangle}}\right].$ (10)
Here, $ij$ are the quark indices, $\theta$ is the complex scaling angle, and $\Psi_{\mathrm{nA}}$ is the decomposed nonantisymmetric wave function, where $Q_{1}\bar{Q}_{3}$ and $Q_{2}\bar{Q}_{4}$ form two color singlets Wu:2024zbx . In the molecular scheme, $r_{13}$ and $r_{24}$ describe the sizes of constituent mesons, while $r_{12},r_{14},r_{23},$ and $r_{34}$ reflect the inter-meson distance. In the compact tetraquark scheme, the rms radii defined by Eq. (10) are consistent with the conventional definition. More discussions on the rms radii can be found in Ref. Wu:2024zbx .
@@ 3.1 Fully heavy tetraquark
We study the S-wave fully charmed tetraquark systems using three different potential models, including the AL1, AP1 and BGS model. We find a series of resonant states, whose energies are listed in Table 2 Wu:2024euj . The three models give qualitatively consistent energy spectra: Most resonances exist in all three models with similar masses and widths. Moreover, the tetraquark spectra in systems with different quantum numbers $J^{PC}$ demonstrates analogous characteristics. A lower state at $M\approx 7$ GeV with width $\Gamma\approx 75$ MeV, and a higher state at $M\approx 7.2$ GeV with width $\Gamma\approx 50$ MeV are found in the $0^{++},1^{+-},2^{++}$ systems. The lower $0^{++}$ and $2^{++}$ states are good candidates for the experimental $X(6900)$ state LHCb:2020bwg , ATLAS:2023bft , CMS:2023owd , while the higher $0^{++}$ and $2^{++}$ states are good candidates for the experimental $X(7200)$ state ATLAS:2023bft , CMS:2023owd . Both the masses and widths agree with the experimental values within tens of MeV. The similar pattern of spectra with different $J^{PC}$ is expected, as the spin splitting term in Eq. (2) is suppressed by the heavy quark mass. However, recent analyses from the CMS collaborations suggest that the experimental states have $J^{PC}=2^{++}$ CMS:2025fpt . Further studies are needed to better understand and resolve the discrepancy.
Table 2: The complex energies $E=M-i\Gamma/2$ (in $\mathrm{MeV}$) of the fully charmed tetraquark resonances from three potential models. The ”?” indicates that potential signs of resonant states are observed but the pole positions cannot be obtained accurately in the present calculations. The ”-” suggests that no corresponding resonance is found. The last column shows possible experimental correspondences of the theoretical states.
$J^{PC}$ AL1 AP1 BGS EXP
$0^{++}$ 698035i6980-35i 697836i6978-36i 703036i7030-36i $X(6900)$
@@
As the spectra from the three models are similar, we choose the results from the AP1 model to study the resonant structures. The color proportions and the rms radii are shown in Table 3. From the rms radii, we find that all fully charmed tetraquark resonances exhibit a compact tetraquark configuration, with four (anti)quarks confined within the typical range of color confinement $\Lambda_{\text{Q}CD}^{-1}\sim 1$ fm. This is expected as the interactions between heavy (anti)quarks should be dominated by short-range gluon exchange. The candidates for $X(6900)$ are dominant by the $\chi_{\bar{3}_{c}\otimes 3_{c}}$ configuration $(\sim 90\%)$, while the candidates for $X(7200)$ show strong mixing effects between $\chi_{\bar{3}_{c}\otimes 3_{c}}$ and $\chi_{6_{c}\otimes\bar{6}_{c}}$.
Table 3: The proportions of different color configurations and the rms radii (in fm) of the $cc\bar{c}\bar{c}$ resonant states in the AP1 potential. The last column shows the spatial configurations of the states, where C. represents the compact tetraquark.
$J^{PC}$ $M-i\Gamma/2$ $\chi_{\bar{3}_{c}\otimes 3_{c}}$ $\chi_{6_{c}\otimes\bar{6}_{c}}$ $r_{c_{1}\bar{c}_{3}}$ $r_{c_{2}\bar{c}_{4}}$ $r_{c_{1}\bar{c}_{4}}$$r_{c_{2}\bar{c}_{3}}$ $r_{c_{1}c_{2}}$$r_{\bar{c}_{3}\bar{c}_{4}}$ Configurations
$0^{++}$ 697836i6978-36i $86\%$ $14\%$ $0.81$ $0.81$ $0.86$ $0.66$ C.
70491i7049-1i $37\%$ $63\%$ $0.70$ $0.70$ $0.82$ $0.75$ C.
716719i7167-19i $46\%$ $54\%$ $0.91$ $0.91$ $0.90$ $0.67$ C.
$1^{+-}$ 69320.5i6932-0.5i $65\%$ $35\%$ $0.66$ $0.66$ $0.73$ $0.63$ C.
699835i6998-35i $88\%$ $12\%$ $0.79$ $0.80$ $0.77$ $0.59$ C.
718127i7181-27i $44\%$ $56\%$ $0.91$ $0.93$ $0.87$ $0.61$ C.
$2^{++}$ 701739i7017-39i $90\%$ $10\%$ $0.79$ $0.79$ $0.71$ $0.56$ C.
71144i7114-4i $69\%$ $31\%$ $0.92$ $0.92$ $0.65$ $0.55$ C.
720429i7204-29i $57\%$ $43\%$ $0.94$ $0.94$ $0.85$ $0.62$ C.
727612i7276-12i $73\%$ $27\%$ $0.86$ $0.86$ $1.04$ $0.93$ C.
We further investigate the S-wave fully bottomed tetraquark Wu:2024euj and fully heavy tetraquark with different flavors Wu:2024hrv within a unified framework. A series of resonant states are predicted in $(19.7,20.0)$ GeV for the $bb\bar{b}\bar{b}$ system and $(13.2,13.5)$ GeV for the $bc\bar{b}\bar{c}$ system, which may be searched for in future experiments. All fully heavy tetraquark resonances are found to be compact tetraquark states. The fully strange tetraquarks, as cousins of the fully heavy tetraquarks, are also studied in Ref. Ma:2024vsi .
@@ 3.2 Doubly heavy tetraquark
As the three potential models are found to give similar results in the fully charmed tetraquark systems, we only choose the AL1 model to investigate the S-wave doubly heavy tetraquark systems $QQ^{(\prime)}\bar{q}\bar{q}$ ($Q^{\prime}=b,c$ and q=u,dq=u,dWu:2024zbx . We observe rich spatial configurations of various doubly heavy tetraquark states by calculating their rms radii. In addition to the general classification of tetraquarks into meson molecules and compact tetraquarks, we further categorize the compact tetraquark configuration into three subtypes, as illustrated in Fig. 2. In compact even tetraquark, the four (anti)quarks are evenly distributed, where their relative distances are of similar size. The two light antiquarks are shared by two heavy quarks, resembling the hydrogen atom, where two electrons are shared by two protons. In compact diquark-antidiquark tetraquark, diquark and antidiquark clusters are formed within the compact structure. In compact diquark-centered tetraquark, two heavy quarks form a compact diquark core, whereas two light antiquarks orbit around the diquark. It corresponds to the QCD helium atom, where two electrons orbit around the helium nucleus. These classifications unravel different internal structures and binding mechanisms of various tetraquark states.
@@
Table 4: The complex energies $E=M-i\Gamma/2$ (in MeV), proportions of color configurations and spatial configurations of the doubly heavy tetraquark bound and resonant states. The binding energies $\Delta E$ of the bound states are shown in the fourth column. The spatial configurations are shown in the last column, where C.E., C.DA., C.DC. and M. respectively indicate compact even tetraquark, compact diquark-antidiquark tetraquark, compact diquark-centered tetraquark and meson molecule. The constituent mesons are shown in the parentheses. The rms radii of the tetraquarks can be found in Ref. Wu:2024zbx . The ”?” indicates that the rms radii are numerically unstable in the complex scaling transformation and the configuration of the state is undetermined.
@@ 3.3 Singly heavy tetraquark
We further study the S-wave charm-strange tetraquark ($cs\bar{q}\bar{q}$) and bottom-strange tetraquark ($bs\bar{q}\bar{q}$) using the AL1 model Chen:2023syh . Table 5 lists the low-lying spectra of the two systems. We identify a series of molecular bound and resonant states, including two scalar bound states ($D\bar{K}$ and $\bar{B}\bar{K}$), two scalar resonances ($D^{*}\bar{K}^{*}$ and $\bar{B}^{*}\bar{K}^{*}$), and one vector bound state ($\bar{B}^{*}\bar{K}$). The $D^{*}\bar{K}^{*}$ molecular resonance stands as a great candidate for the experimental $T^{*}_{\bar{c}\bar{s}0}(2870)$, whose mass and width are $2872\pm 16$ MeV and $67\pm 24$ MeV ParticleDataGroup:2024cfk . The theoretical width is expected to be smaller than the experimental value due to the neglection of the width of $\bar{K}^{*}$ in the calculations.